Fermi National Laboratory
 

Operation Summary for September 27, 2010 - September 29, 2010

 

Monday Sept. 27

The day shift began with Operations monitoring the Tevatron (TeV) store 8127 (176.03E30) and the Recycler stash (76.7E10), with the Antiproton Source (Pbar) stacking (19.0E10), with the Linac, Booster, and the Main Injector (MI) operating normally, with NuMI off for target work, with the MiniBooNE and the Meson FTFB T992 and T995 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, Linac, Booster, and Preaccelerator personnel performed maintenance. A Linac RF station (LRF3) and Booster RF stations (BRF9 and 19) tripped off several times during shift.

During the evening shift, Operations disabled beam to T992 and T995 at 6 PM. Operations terminated store 8127 at 9:37 PM. Operators prepared the CDF and D0 collision halls for accesses. A kicker in the Recycler beam line faulted on a prefire indication causing the loss of 115E10 of beam from the stash. No cause was found. The kicker was reset and operated normally thereafter. CDF and D0 were out of access by the end of shift.

Tuesday Sept. 28

The midnight shift began with Operations monitoring the stash (295.8E10), with Pbar stacking (10.8E10), and with MiniBooNE taking beam.

During the midnight shift, Operations put the TeV into HEP shot setup at 1:06 AM and established store 8130 at 2:05 AM with an average initial luminosity (AIL) of 289.18E30. Operations established beam to T992 and T995 at 4:22 AM.

During the day shift, Operations terminated store 8130 at 3:55 PM. EE Support personnel worked on a TeV HOPS power supply. Operations reported several minor downtimes due to Linac and Booster RF systems.

During the evening shift, the TeV power supply work finished. Operations put the TeV into shot setup at 4:33 PM and established store 8131 at 5:31 PM with an AIL of 321.99E30.

Wednesday Sept. 29

The midnight shift began with Operations monitoring store 8131 (126.2E30) and the stash (190.8E10), with Pbar stacking (28.8E10), and with MiniBooNE taking beam.

During the midnight shift, Operations disabled the MiniBooNE beam at 4 AM to prepare for the experiment’s day shift access. Operations established beam to T992 and T995 at 4:14 AM. Operations terminated store 8131 at 7:29 AM and put the TeV into shot setup. A Recycler kicker voltage read-back cable failed; EE Support personnel came in and repaired the cable.

The Plans for Wednesday

The plans for today are to stack, stash, and store. The Run Coordinator will allow MI personnel to conduct collimator studies, Recycler personnel to conduct studies, and then after the Neutron Therapy Facility treats its last patient for the day, he will allow the controls personnel to reboot the Time Line Generator.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 09/30/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory